Grade 12th pass9 grade science

Two bodies of different masses m1​ and m2​ are dropped from two different heights div​ and h2​. What is the ratio of time taken by the two to reach the ground?

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
For a freely falling object :
a=g;

Initial velocity=u=0 m/s

g=9.8m/s²

from second equation of motion:S=ut+1/2 at²


for First object we get :


h1= gT1²/2

⇒T1=√2h1/g


For second object we get h2=gT2²/2

⇒T2=√2h2/g
T1/T2 =√h1/h2

But here h1= a and h2 =b
so, T1/T2=√a/b
THANKS